OpenStudy (anonymous):

LCD of 9/10+ 2/5=...either 11/15 or 1 3/10

hero (hero):

1. Find the LCM of 10 and 5

OpenStudy (anonymous):

9/10 +2/5(2/2)= 9/10+4/10=13/10 -> 1 3/10

OpenStudy (anonymous):

MCronin u r the BOMB...thank u for your help

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

really,I thank u so much,I have no idea and this is mt sons math homework,looks like I need to start relearning math all-over LoL thank you,good night

hero (hero):

I wanted to lead you to the answer, not just give it to you.

OpenStudy (anonymous):

thankx HERO